OpenStudy (anonymous):

Triangles ABC and LMN are given such that AB = 24, BC = 20, and AC = 40 LM = 6, MN = 5, and LN = (3a + 1) To be similar, which should be the value of a? 6 5 4 3

OpenStudy (anonymous):

40/(3a+1)=4 4 is the similarity ratio which can be equal to 24/6 or 20/5 40/(3a+1)=4 40=12a+4 36=12a a=36/12 a=3

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

but i dont think they framed it correctly coz 20/10=2 and the ratio was found to be 4 but thats my answer welcome